# Transport properties through alternating borophene and graphene superlattices

### v1

@@ Abstract
The electronic transport properties of two junctions (BGB, GBG) made of borophene (B) and graphene (G) are studied. Using the transfer matrix method with Chebyshev polynomials, we have studied single and multiple barriers in a superlattice configuration. We showed that a single barrier exhibits remarkable tilted transport properties, with perfect transmission observed for both junctions under normal incidence. We found that robust superlattice transmission is maintained for multiple barriers, particularly in the BGB junction. It turns out that by varying the incident energy, many gaps appear in the transmission probability. The number, width, and position of these transmission gaps can be manipulated by adjusting the number of cells, incident angle, and barrier characteristics. For diffuse transport, we observed significant variations in conductance and the Fano factor, highlighting the sensitivity of these junctions to the physical parameters. We showed different behaviors between BGB and GBG junctions, particularly with respect to the response of conductance and Fano factor when barrier height varies. For ballistic transport, we have seen that the minimum conductivity is related to the maximum Fano factor, demonstrating their control under specific conditions of the physical parameters. Analysis of the length ratio (geometric factor) revealed some remarkable patterns, where conductivity and the Fano factor converged to certain values as the ratio approached infinity.
@@ I Introduction
The discovery of graphene, a single layer (SLG) of carbon atoms arranged in a hexagonal lattice, was a milestone in materials science and earned the Nobel Prize in 2010 [1, 2]. The two-dimensional wonder material demonstrated superior electrical, mechanical, and thermal properties, laying the foundations for a wide range of applications [3, 4, 5, 6]. Following in the footsteps of graphene, borophene, a monolayer of boron atoms with unique honeycomb structures, has emerged as a promising nanomaterial [7, 8, 9, 10]. Different phases of bulk and two-dimensional (2D) boron allotropes, such as $\alpha$, $\beta$, and others, have been put forward in theoretical proposals [11, 12, 13, 14, 7]. Although boron typically avoids forming chemical bonds to maintain a stable honeycomb lattice, it is feasible to generate a stable planar structure by integrating honeycomb with triangular units  [15, 16]. This structure, known as $2B:Pmmn$, has two atoms per primitive unit cell and belongs to the orthorhombic crystal system with space group 59 ($Pmmn$). A novel Dirac material known as hydrogenated borophene, or borophane, was predicted [17] to exhibit Dirac properties with a remarkable Fermi velocity nearly twice that of graphene.
@@
Characterized by interesting properties different from those of graphene, borophene has emerged as a fascinating facet in the field of 2D materials. The importance of studying graphene and borophene lies in their different contributions to nanoscale physics and materials science, highlighting the need for further exploration of these materials [18, 19, 20, 21, 22]. By exploring the interfaces between two-dimensional materials, one can discover a realm of intricate interactions and emergent phenomena [23, 24, 25]. The key to unlocking new functionalities lies at these interfaces, where the unique properties of individual materials come together. The integration of graphene and borophene junctions has the potential to provide unprecedented electronic, structural, and transport properties [26, 27]. The study of this interface becomes not only a quest to understand the interaction of graphene and borophene, but also a gateway to exploit their combined potential for innovative applications in nanoscale devices and materials engineering [28, 29, 30]. Notable contributions from various studies highlight the importance of understanding this interface, with research on graphene junctions [31, 32, 33, 34] and borophene [35, 36] serving as crucial steps.
@@
Building on previous advances in understanding the transport properties of graphene (G) and borophene (B) in the presence of various potential barriers, we undertake a similar investigation focused on analyzing the behavior of two junctions, BGB and GBG. Specifically, we study the case where these two junctions encounter a single barrier and extend it to scenarios with multiple barriers, similar to a superlattice. Consequently, we have arrived at the key findings: In the case of a single barrier, our study has uncovered remarkable electronic transport properties. Indeed, regardless of the barrier width or height, perfect transmission is consistently observed at normal incidence in both junctions. However, the BGB junction has a significant effect, showing the influence of the potential barrier, as reducing the barrier height results in a loss of transmission. We observe pronounced scattered transmission at normal incidence as long as the energy and transverse wave vector $k_{y}$ vary. In addition, we show that the single barrier induces resonances associated with the finite size of graphene and borophene. In the scenario involving multiple barriers (superlattice), we observe remarkable transport properties. Notably, the superlattice maintains perfect transparency for the BGB junction even at non-null incident angles, highlighting robust transmission. Resonance peaks, particularly notable in BGB, depict distinctive transport phenomena that are affected by the number and position of the cells. Specifically, as the energy varies, multiple transmission gaps emerge. The number, width, and position of these gaps can be modified by adjusting the number of cells, incident angle, and barrier parameters. In the case of diffuse transport, we find significant variations in conductance and Fano factor, highlighting the sensitivity of BGB and GBG junctions to physical parameters. In particular, we observed different behaviors between BGB and GBG junctions, especially in terms of how the conductance and Fano factor respond to variations in barrier height. In the case of ballistic transport, an intriguing relationship emerged between the minimum conductivity and the maximum Fano factor. This relationship demonstrated the controllability of these factors under certain conditions of the physical parameters. In addition, our analysis of the length ratio, also known as the geometric factor, revealed remarkable patterns. As the ratio approached infinity, the conductivity and the Fano factor converged to certain values, suggesting intriguing properties associated with this geometric configuration.
@@
Figure 1: (Color online) Band dispersion relations around the Dirac point for BGB-junction (a) and GBG-junction (b) are plotted in each region of a rectangular potential barrier of height $V_{0}$ and width $d$. The blue tilted energy and red linear energy dispersions correspond to borophene (B) and graphene (G), respectively. (c) defines the 1D periodic potential of $(\text{BG})^{n}$ (case 1) and $(\text{GB})^{n}$ (case 2) superlattices, where $n$ is the number of cells, $d$ is the barrier width, and $w$ is the well width. We have set the period $L=d+w$.
@@
The paper is organized as follows. In Sec. II, we present a theoretical model describing fermions in borophene and graphene. In Sec. III, we use a mathematical formalism based on the continuity of spinors at interfaces and the transfer method to compute the transmission, conductance, and Fano factor. In Sec. IV, we numerically analyze our results by giving different discussions and comparisons with literature. We summarize our main conclusions in Sec. V.
@@ II Theoretical MODEL
We offer separate presentations for each configuration: borophene-graphene-borophene (BGB) and graphene-borophene-graphene (GBG), as depicted in Fig. 1 (a, b). To study the transport properties in both junctions and provide comparisons, we consider the periodic potential illustrated in Fig. 1 (c) and given by
@@ II.1 Borophene
We initiate our study with the low-energy continuum Hamiltonian for a monolayer borophene in region $j$th, which describes a tilted anisotropic Dirac cone in 2D using three parameters [37, 38, 39, 40]
@@ II.2 Graphene
with the complex number $g_{j}=\frac{1}{g_{j}^{*}}=\text{sgn}(\varepsilon_{\text{G},j})\frac{q_{x,j}\pm ik_{y}}{\varepsilon_{\text{G},j}}$, $E_{\text{G},j}-V_{j}=\varepsilon_{\text{G},j}$, and the wave vector $q_{x,j}$ is defined
@@ III TRANSPORT quantities
By ensuring the continuity of the wave function across the various boundaries between barriers and wells, we can establish the transfer matrix for the periodic structures BGB and GBG, each composed of $n$ unit cells, respectively,
@@
$\cos(K_{x}L)$ $=\alpha\left[(g_{2}-f_{1}^{-})(1+f_{1}^{+}g_{2})\cos(dk_{x,1}+wq_{x,2})+(f_{1}^{+}-g_{2})(1+f_{1}^{-}g_{2})\cos(dk_{x,1}-wq_{x,2})\right],$ (16)
$\cos(K_{x}L)$ $=\beta\left[(g_{1}-f_{2}^{-})(1+f_{2}^{+}g_{1})\cos(dq_{x,1}+wk_{x,2})+(f_{2}^{+}-g_{1})(1+f_{2}^{-}g_{1})\cos(dq_{x,1}-wk_{x,2})\right],$ (17)
where we have set the parameters $\alpha=\frac{1}{(1+g_{2}^{2})(f_{1}^{+}-f_{1}^{-})}$ and $\beta=(1+g_{1}^{2})(f_{2}^{+}-f_{2}^{-})$, $K_{x}$ is the Bloch wave vector.
In Fig. 2, we illustrate the band structures for BGB (Eq. (16)) and GBG (Eq. (17)) under appropriate conditions of the physical parameters. This sheds light on the impact of the incident angle on the mini-zone boundary. In particular, enlarged depictions of the band structures near the wave vector $\pi/L$ are presented for BGB and GBG junctions by choosing the barrier height between the two layers $V_{0}=0$ (a) and $V_{0}=50$ meV (b, c), where $d=w=20$ nm. In Fig. 2(a) for $V_{0}=0$, a notable observation arises in the band structures of BGB and GBG junctions. Regardless of the values taken by the incident angle $\theta$, both BGB and GBG band structures remain unchanged with no gap opening, unlike the findings observed in graphene superlattices [42]. This observation underscores the crucial role of barrier height in altering the band structure of graphene-based heterostructures. Subsequently, in Figs. 2(b, c), we observe that the energy separation around ±$\pm\pi/L$ depends on both the wave vector and incident angle. Notably, at $\theta=0^{\circ}$, there is no gap opening at the center of the mini-zone boundary for both the two junctions. However, at $\theta=10^{\circ}$ and $20^{\circ}$, the original Dirac cones have upward shifts, and the gap exhibits sensitivity to the incident angle.
@@
where $U_{n}(x)=\frac{\sin[(n+1)x]}{\sin x}$ represent Chebychev polynomials of the second kind, $\omega_{\text{BGB/GBG}}=\arccos\left[\frac{1}{2}\mathrm{tr}(M_{1,\text{BGB/GBG}})\right]$, and $\Omega_{\text{BGB/GBG},ij}$ are the $(i,j)$ elements of the single period matrices Eqs. (14, 15). This process results in acquiring the transmission coefficient associated with the $n$ barrier, as
@@
where $\theta=\arccos\frac{Q_{x}}{E_{F}}$, with $Q_{x}=k_{x}$ for borophene and $Q_{x}=q_{x}$ for graphene. Here, $G_{0}=2e^{2}E_{F}D/(\pi\hbar)$ is the conductance unit, and $D$ is the sample size in the $y$-direction. Finally, the Fano factor is given by
@@ IV.1 Single barrier (one cell)
In Fig. 3, we present the transmission probabilities as a function of the incident angle $\theta$ in the absence of potential barrier $V_{0}=0$ for two barrier widths: (a) $d=50$ nm and (b) $d=100$ nm.  Remarkably, perfect transmission is observed under normal incidence for both BGB and GBG junctions, regardless of the particular value of $d$, resulting in signature of Klein tunneling. i.e., $T=1$. This robust behavior highlights the exceptional transport characteristics of these structures, especially concerning the normal incident angle and $V_{0}=0$. This effect remains consistent even when the potential barrier is non-zero [41]. In contrast to the findings in [41], the transmissions exhibit asymmetry for both junctions at non-normal incidence. Furthermore, perfect transmission occurs over a significant interval of incident angles associated with each junction. This phenomenon of perfect tunneling can be explained by invoking the principle of pseudospin conservation between the energy states of graphene and borophene.
@@
Figure 3: (Color online) Angular dependence of transmission probability for BGB and GBG junctions with $E=82,49$ meV, $V_{0}=0$, and two barrier widths: (a) $d=50$ nm, (b) $d=100$ nm. Red curves represent BGB, and blue curves represent GBG.
Figure 4: (Color online) Contour plot of the transmission probability as a function of the energy and transverse wave vector for $V_{0}=0$ and  $d=100$ nmm with BGB (a) and GBG (b). The dashed black lines represent the band energy for graphene.
@@
While the previous analysis focused on junction behaviors independent of potential barriers, our investigation now extends to evaluating the consequences under the influence of various potential barriers. In Fig. 5, we show the angular dependence of transmission probability (a) BGB and (b) GBG, but now in the presence of the potential barrier $V_{0}\not=0$. The discovery shows distinct patterns in electronic transport when Dirac fermions are normally incident at these junctions. Particularly, we observe a perfect transmission at normal incidence, indicating distinct characteristics of Dirac fermions within these heterostructures. For the BGB junction with $V_{0}=285$ meV, the results align with those obtained in [41] by studying monolayer graphene. However, a significant shift occurs when the barrier height is reduced to $200$ meV, resulting in vanishing transmission.  The asymmetry in transmission at non-normal incidence for the two junctions is consistently achieved, mirroring the previous scenario where $V_{0}=0$. Furthermore, these resonances emerge even when the incident energy is below the barrier height, a phenomenon grounded in the conservation of pseudospin, directly associated with the Klein paradox in quantum electrodynamics [41]. The positions and numbers of these resonances primarily hinge on the barrier height and the specific junction.
A contour plot of transmission probabilities as a function of energy and transverse wave vector is shown in Fig. 6 for $V_{0}=0.3$ eV and $d=100$ nm, such that BGB (a) and GBG (b). This figure shows a $(k_{y},E)$ plane, identifying distinct quantifiable regions that signify different modes both inside and outside the junctions. Remarkably, there is an asymmetry in the transmissions at $k_{y}=0$, where $d$ significantly affects the scattered transmission due to the absence of valley equivalence. In the presence of the potential barrier in the central region induces resonances in the transmission probabilities for energies in the range $E<V_{0}$. This phenomenon is associated with the finite size of graphene and borophene, as well as the presence of charge carriers with different chirality. Importantly, there is a noticeable difference between GBG and BGB. This can be attributed to the lack of valley equivalence, where carriers scattering from borophene to graphene differ from those scattering in the opposite direction. It is noteworthy that at $k_{y}=0$, Klein tunneling is observed in graphene [41] through a potential barrier, even at low applied potentials of $E<V_{0}$. Also, it emerges here for both BGB and GBG junctions, showing that the barrier consistently remains perfectly transparent.

### v2

@@ Abstract
The electronic transport properties of two junctions (BGB, GBG) made of borophene (B) and graphene (G) are investigated. Using the transfer matrix method with Chebyshev polynomials, we have studied single and multiple barriers in a superlattice configuration. We showed that a single barrier exhibits remarkable tilted transport properties, with perfect transmission observed for both junctions under normal incidence. We found that robust superlattice transmission is maintained for multiple barriers, particularly in the BGB junction. It turns out that by varying the incident energy, many gaps appear in the transmission probability. The number, width, and position of these transmission gaps can be manipulated by adjusting the number of cells, incident angle, and barrier characteristics. For diffuse transport, we observed considerable variations in transmission probability, conductance and the Fano factor, highlighting the sensitivity of these junctions to the physical parameters. We showed different behaviors between BGB and GBG junctions, particularly with respect to the response of conductance and Fano factor when barrier height varies. For ballistic transport, we have seen that the minimum scaled conductance is related to the maximum Fano factor, demonstrating their control under specific conditions of the physical parameters. Analysis of the length ratio (geometric factor) revealed some remarkable patterns, where scaled conductance and the Fano factor converged to certain values as the ratio approached infinity.
@@ I Introduction
The discovery of graphene, a single layer of carbon atoms arranged in a hexagonal lattice, was a milestone in materials science and earned the Nobel Prize in 2010 [1, 2]. The two-dimensional (2D) wonder material demonstrated superior electrical, mechanical, and thermal properties, laying the foundations for a wide range of applications [3, 4, 5, 6]. Following in the footsteps of graphene, borophene, a monolayer of boron atoms with unique honeycomb structures, has emerged as a promising nanomaterial [7, 8, 9, 10]. Different phases of bulk and two-dimensional (2D) boron allotropes, such as $\alpha$, $\beta$, and others, have been put forward in theoretical proposals [11, 12, 13, 14, 7]. Although boron typically avoids forming chemical bonds to maintain a stable honeycomb lattice, it is feasible to generate a stable planar structure by integrating honeycomb with triangular units [15, 16]. This structure, known as $2B:Pmmn$, has two atoms per primitive unit cell and belongs to the orthorhombic crystal system with space group 59 ($Pmmn$). A novel Dirac material known as hydrogenated borophene, or borophane, was predicted [17] to exhibit Dirac properties with a remarkable Fermi velocity nearly twice that of graphene.
@@
Characterized by interesting properties different from those of graphene, borophene has emerged as a fascinating facet in the field of 2D materials. The importance of studying graphene and borophene lies in their different contributions to nanoscale physics and materials science, highlighting the need for further exploration of these materials [18, 19, 20, 21, 22]. By exploring the interfaces between 2D materials, one can discover a realm of intricate interactions and emergent phenomena [23, 24, 25]. The key to unlocking new functionalities lies at these interfaces, where the unique properties of individual materials come together. The integration of graphene and borophene junctions has the potential to provide unprecedented electronic, structural, and transport properties [26, 27]. The study of this interface becomes not only a quest to understand the interaction of graphene and borophene, but also a gateway to exploit their combined potential for innovative applications in nanoscale devices and materials engineering [28, 29, 30]. Notable contributions from various studies highlight the importance of understanding this interface, with research on graphene [31, 32, 33, 34] and borophene [35, 36] junctions serving as crucial steps.
@@
Building on previous advances in understanding the transport properties of graphene (G) and borophene (B) in the presence of various potential barriers, we undertake a similar investigation focused on analyzing the behavior of two junctions, BGB and GBG. Specifically, we study the case where these two junctions encounter a single barrier and extend it to scenarios with multiple barriers, similar to a superlattice. Consequently, we have arrived at the key findings: In the case of a single barrier, our study has uncovered remarkable electronic transport properties. Indeed, regardless of the barrier width or height, perfect transmission is consistently observed at normal incidence in both junctions. However, the BGB junction has a predominant effect, showing the influence of the potential barrier, as reducing the barrier height results in a loss of transmission. We observe pronounced scattered transmission at normal incidence as long as the energy and transverse wave vector $k_{y}$ vary. In addition, we show that the single barrier induces resonances associated with the finite size of graphene and borophene. In the scenario involving multiple barriers (superlattice), we observe remarkable transport properties. Notably, the superlattice maintains perfect transparency for the BGB junction even at non-null incident angles, highlighting robust transmission. Resonance peaks, particularly notable in BGB, depict distinctive transport phenomena that are affected by the number and position of the cells. Specifically, as the energy varies, multiple transmission gaps emerge. The number, width, and position of these gaps can be modified by adjusting the number of cells, incident angle, and barrier characteristics. In the case of diffuse transport, we find considerable variations in transmission probability, conductance and Fano factor, highlighting the sensitivity of BGB and GBG junctions to physical parameters. In particular, we observed different behaviors between BGB and GBG junctions, especially in terms of how the conductance and Fano factor respond to variations in barrier height. In the case of ballistic transport, an intriguing relationship emerged between the minimum scaled conductance and the maximum Fano factor. This relationship demonstrated the controllability of these factors under certain conditions of the physical parameters. In addition, our analysis of the length ratio, also known as the geometric factor, revealed remarkable patterns. As the ratio approached infinity, the scaled conductance and the Fano factor converged to certain values, suggesting intriguing properties associated with this geometric configuration.
@@
Figure 1: (Color online) (a) Band dispersion relations around the Dirac point for BGB-junction and GBG-junction are plotted in each region of a rectangular potential barrier of height $V_{0}$ and width $d$. The blue tilted energy and red linear energy dispersions correspond to borophene (B) and graphene (G), respectively. (b) defines the 1D periodic potential of $(\text{BG})^{n}$ (case 1) and $(\text{GB})^{n}$ (case 2) superlattices, where $n$ is the number of cells, $d$ is the barrier width, and $w$ is the well width. We have set the period to $L=d+w$.
@@
The paper is organized as follows. In Sec. II, we present a theoretical model describing fermions in borophene and graphene. In Sec. III, we use a mathematical formalism based on the continuity of spinors at interfaces and the transfer method to compute the transmission, conductance, and Fano factor. In Sec. IV, we numerically analyze our results by offering different discussions and comparisons with literature. We summarize our main conclusions in Sec. V.
@@ II Theoretical MODEL
In Fig. 1, we illustrate the two configurations: borophene-graphene-borophene (BGB) and graphene-borophene-graphene (GBG). According to Fig. 1(a), one sees that both borophene and graphene exhibit linear dispersions typical of Dirac cones. These are visible within a small energy range known as the low-energy approximation [37]. To study the transport properties in both junctions and provide comparisons, we consider the periodic potential illustrated in Fig. 1 (b) and given by
@@ II.1 Borophene
We initiate our study with the low-energy continuum Hamiltonian for a monolayer borophene in $j$th region, which describes a tilted anisotropic Dirac cone in 2D using three parameters [38, 39, 40, 41]
@@ II.2 Graphene
with the complex number $g_{j}=\frac{1}{g_{j}^{*}}=\text{sgn}(\varepsilon_{\text{G},j})\frac{q_{x,j}\pm ik_{y}}{\varepsilon_{\text{G},j}}$, $E_{\text{G},j}-V_{j}=\varepsilon_{\text{G},j}$, and the wave vector $q_{x,j}$ is defined by
@@ III TRANSPORT quantities
By ensuring the continuity of the wave function across the various boundaries between barriers and wells, we can establish the transfer matrix for the periodic structures BGB and GBG [43, 44, 45], each composed of $n$ unit cells, respectively,
@@
$\cos(K_{x}L)$ $=\alpha\left[\left(g_{2}-f_{1}^{-}\right)\left(1+f_{1}^{+}g_{2}\right)\cos(dk_{x,1}+wq_{x,2})+\left(f_{1}^{+}-g_{2}\right)\left(1+f_{1}^{-}g_{2}\right)\cos(dk_{x,1}-wq_{x,2})\right],$ (16)
$\cos(K_{x}L)$ $=\beta\left[\left(g_{1}-f_{2}^{-}\right)\left(1+f_{2}^{+}g_{1}\right)\cos(dq_{x,1}+wk_{x,2})+\left(f_{2}^{+}-g_{1}\right)\left(1+f_{2}^{-}g_{1}\right)\cos(dq_{x,1}-wk_{x,2})\right],$ (17)
where we have set the parameters $\alpha=\frac{1}{\left(1+g_{2}^{2}\right)\left(f_{1}^{+}-f_{1}^{-}\right)}$ and $\beta=\left(1+g_{1}^{2}\right)\left(f_{2}^{+}-f_{2}^{-}\right)$, $K_{x}$ is the Bloch wave vector.
In Fig. 2, we illustrate the band structures for BGB (Eq. (16)) and GBG (Eq. (17)) under appropriate conditions of the physical parameters. This sheds light on the impact of the incident angle on the mini-zone boundary. In particular, enlarged depictions of the band structures near the wave vector $\pi/L$ are presented for BGB and GBG junctions by choosing the barrier height between the two layers $V_{0}=0$ (a) and $V_{0}=50$ meV (b, c), where $d=w=20$ nm. In Fig. 2(a) for $V_{0}=0$, a notable observation arises in the band structures of BGB and GBG junctions. Regardless of the values taken by the incident angle $\theta$, both BGB and GBG band structures remain unchanged with no gap opening, unlike the findings observed in graphene superlattices [46]. This observation underscores the critical role of barrier height in scenarios involving non-zero potential barriers, which will be discussed in the following sections. Subsequently, in Figs. 2(b, c), we observe that the energy separation around ±$\pm\pi/L$ depends on both the wave vector and incident angle. Notably, at $\theta=0^{\circ}$, there is no gap opening at the center of the mini-zone boundary for both the two junctions. However, at $\theta=10^{\circ}$ and $20^{\circ}$, the original Dirac cones have upward shifts, and the gap exhibits sensitivity to the incident angle.
@@
where $U_{n}(x)=\frac{\sin[(n+1)x]}{\sin x}$ represent Chebychev polynomials of the second kind, $\omega_{\text{BGB/GBG}}=\arccos\left[\frac{1}{2}\mathrm{tr}\left(M_{1,\text{BGB/GBG}}\right)\right]$, and $\Omega_{\text{BGB/GBG},ij}$ are the $(i,j)$ elements of the single period matrices Eqs. (14, 15). This process results in acquiring the transmission coefficient associated with the $n$ barrier, as
@@
where $\theta=\arccos\frac{Q_{x}}{E_{F}}$, with $Q_{x}=k_{x}$ for borophene and $Q_{x}=q_{x}$ for graphene. Here, $G_{0}=2e^{2}E_{F}D/(\pi\hbar)$ is the conductance unit, and $D$ is the sample size in the $y$-direction. Finally, the Fano factor is given by [50, 51]
@@ IV.1 Single barrier (one cell)
In Fig. 3, we present the transmission probabilities as a function of the incident angle $\theta$ in the absence of potential barrier ($V_{0}=0$) for two barrier widths: (a) $d=50$ nm and (b) $d=100$ nm. Remarkably, perfect transmission is observed under normal incidence for both BGB and GBG junctions, regardless of the particular value of $d$, resulting in signature of Klein tunneling. i.e., $T=1$. This robust behavior highlights the exceptional transport properties of these structures, especially concerning the normal incident angle $(\theta=0)$ and $V_{0}=0$. This effect remains consistent even when the potential barrier is non-zero. In contrast to the findings in [42], the transmissions exhibit asymmetry for both junctions at non-normal incidence. Furthermore, perfect transmission occurs over a wide range of incident angles associated with each junction. This phenomenon of perfect tunneling can be explained by invoking the principle of pseudospin conservation between the energy states of graphene and borophene.
@@
Figure 3: (Color online) Angular dependence of transmission probability for BGB and GBG junctions with $E=82.49$ meV, $V_{0}=0$, and two barrier widths: (a) $d=50$ nm, (b) $d=100$ nm. Red curves represent BGB, and blue curves represent GBG.
Figure 4: (Color online) Contour plot of the transmission probability as a function of the energy and transverse wave vector for $V_{0}=0$ and $d=100$ nm with BGB (a) and GBG (b). The dashed black lines represent the band energy for graphene.
@@
While the previous analysis focused on junction behaviors independent of potential barriers, our investigation now extends to evaluating the consequences under the influence of various potential barriers. In Fig. 5, we show the angular dependence of transmission probability (a) BGB and (b) GBG, but now in the presence of the potential barrier ($V_{0}\not=0$). The discovery shows distinct patterns in electronic transport when Dirac fermions are normally incident at these junctions. Particularly, we observe a perfect transmission at normal incidence, indicating distinct characteristics of Dirac fermions within these heterostructures. For the BGB junction with $V_{0}=285$ meV, the results align with those obtained in [42] by studying monolayer graphene. However, a large shift occurs when the barrier height is reduced to $200$ meV, resulting in vanishing transmission. The asymmetry in transmission at non-normal incidence for the two junctions is consistently achieved, mirroring the previous scenario where $V_{0}=0$. Furthermore, these resonances emerge even when the incident energy is below the barrier height, a phenomenon grounded in the conservation of pseudospin, directly associated with the Klein paradox in quantum electrodynamics [42]. The positions and numbers of these resonances primarily hinge on the barrier height and the specific junction.
A contour plot of transmission probabilities as a function of energy and transverse wave vector is shown in Fig. 6 for $V_{0}=0.3$ eV and $d=100$ nm, such that BGB (a) and GBG (b). This figure shows a $(k_{y},E)$ plane, identifying distinct quantifiable regions that signify different modes both inside and outside the junctions. Remarkably, there is an asymmetry in the transmissions at $k_{y}=0$, which is due to the absence of valley equivalence [52]. In the presence of the potential barrier in the central region induces resonances in the transmission probabilities for energies in the range $E<V_{0}$. This phenomenon is associated with the finite size of graphene and borophene, as well as the presence of charge carriers with different chirality. Importantly, there is a noticeable difference between GBG and BGB. This can be attributed to the lack of valley equivalence, where carriers scattering from borophene to graphene differ from those scattering in the opposite direction. It is noteworthy that at $k_{y}=0$, Klein tunneling is observed in graphene [42] through a potential barrier, even at low applied potentials of $E<V_{0}$. Also, it emerges here for both BGB and GBG junctions, showing that the barrier consistently remains perfectly transparent.